If the consumption function of a given individual is given C= 44+0.86Yd and the individual’s disposable income for a specific period was birr 3600, 


1
Expert's answer
2022-01-11T14:36:04-0500

If the consumption function of a given individual is given C= 44+0.86Yd and the individual’s disposable income for a specific period was birr 3600, then his consumption is:

"C= 44+0.86\u00d73600 = 3140."